# What’s Inside Your Diffusion Model? A Score-Based Riemannian Metric to Explore the Data Manifold

### v2

@@ Abstract
Recent advances in diffusion models have demonstrated their remarkable ability to capture complex image distributions, but the geometric properties of the learned data manifold remain poorly understood. We address this gap by introducing a score-based Riemannian metric that leverages the Stein score function from diffusion models to characterize the intrinsic geometry of the data manifold without requiring explicit parameterization. Our approach defines a metric tensor in the ambient space that stretches distances perpendicular to the manifold while preserving them along tangential directions, effectively creating a geometry where geodesics naturally follow the manifold’s contours. We develop efficient algorithms for computing these geodesics and demonstrate their utility for both interpolation between data points and extrapolation beyond the observed data distribution. Through experiments on synthetic data with known geometry, Rotated MNIST, and complex natural images via Stable Diffusion, we show that our score-based geodesics capture meaningful transformations that respect the underlying data distribution. Our method consistently outperforms baseline approaches on perceptual metrics (LPIPS) and distribution-level metrics (FID, KID), producing smoother, more realistic image transitions. These results reveal the implicit geometric structure learned by diffusion models and provide a principled way to navigate the manifold of natural images through the lens of Riemannian geometry.
@@ 1 Introduction
The geometry of natural images can be studied by viewing them as points on a manifold in high-dimensional ambient – pixel – space. This perspective, commonly known as the manifold hypothesis [1, 2], suggests that despite the vast dimensionality of pixel space, natural images concentrate near a much lower-dimensional structure. Understanding this geometric structure is crucial for numerous applications including image synthesis, manipulation, and analysis [3, 4], visual perception [5, 6], and representation learning [7, 8, 9].
In this work, we propose a data-dependent metric derived from the Stein score function, which can be obtained from a diffusion model trained on a given set of images. Recent advances in diffusion models have demonstrated remarkable capabilities in capturing complex image distributions [10, 11], and several approaches have explored data-dependent metrics to uncover the geometrical properties of image manifolds [12, 13, 14]. However, these previous methods often require explicit manifold parameterization, are computationally intractable for high-dimensional data, or fail to capture the intrinsic geometric structure that reflects perceptual relationships in the data.
Our approach defines a Riemannian metric in the ambient – pixel – space, leveraging score functions from diffusion models to create a geometry that stretches perpendicular to the data manifold while preserving distances along it. This metric allows us to compute distances and geodesic paths between images directly in pixel space, capturing meaningful relationships that respect the underlying data distribution. Unlike the Euclidean distance, which treats all directions in pixel space equally, our score-induced metric accounts for the anisotropic nature of the image manifold, heavily penalizing movement in directions orthogonal to the manifold while preserving natural movement along tangential directions. By generating images along these geodesics, by exploiting a diffusion model, we can visualize and validate these transformations, providing unique insights into the structure of the learned image manifold.
The main contributions of our work are:
  • A novel score-based Riemannian metric that captures the geometry of the data manifold without requiring explicit parameterization
  • Efficient algorithms for computing geodesics and manifold extrapolation in high-dimensional image space
  • Empirical validation on both synthetic data with known geometry and complex image data
@@ 2.1 The Manifold Hypothesis for Natural Images
The manifold hypothesis posits that high-dimensional data, such as natural images, approximately lies on a low-dimensional manifold embedded in the ambient space $\mathbb{R}^{N}$ [1, 2]. In this framework, images are represented as points in a $N$-dimensional pixel space, where each axis represents a pixel value. Despite the vast dimensionality of this space – often millions of dimensions for modern high-resolution images –, natural images occupy only a tiny fraction of it due to strong correlations and constraints that restrict the set of plausible pixel configurations [15].
Each set of images can be characterized by its own probability density function $p(\boldsymbol{x})$ defined over the pixel space $\mathbb{R}^{N}$. The key insight is that $p(\boldsymbol{x})$ concentrates its mass on a much lower-dimensional structure than the ambient dimension $N$, referred to as the support of the data distribution. Under the manifold hypothesis, this support approximates a lower-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$ (Figure 1 A). To capture the geometry of this implicit manifold, we employ data-induced metrics - Riemannian metrics defined in the ambient space that adapt locally according to the underlying probability density [16]. Our approach explores how ambient space deformation enables geodesics that naturally follow the manifold’s contours (Figure 1 B,C).
Figure 1: (A) Data points $\mathbf{x}$ sampled from a probability distribution $p(\mathbf{x})$ (left) concentrate on a lower-dimensional manifold $\mathcal{M}$ (right). (B) Linear interpolation (black dashed) versus geodesic interpolation (red curve) between MNIST digits. Geodesics follow the manifold surface, producing valid digit transitions, while linear paths yield superpositions. (C) Geometric deformation: data manifold $\mathcal{M}$ embedded in $\mathbb{R}^{N}$ (left) transforms flat Euclidean space into curved metric space ~$\tilde{\mathbb{R}}^{N}$ (middle, right). In this deformed space, geodesics of $\mathcal{M}$ can be computed directly as geodesics of $\tilde{\mathbb{R}^{N}}$.
@@ 2.2 Diffusion Models & Stein Score
Diffusion models have emerged as powerful generative approaches that gradually transform noise into structured data through an iterative denoising process [17, 10]. These models define a forward process $q$ that progressively adds Gaussian noise to data points $\mathbf{x}_{0}\sim p(\mathbf{x})$ according to a predefined schedule, creating a sequence of increasingly noisy versions $\mathbf{x}_{t}$ with $t$ representing the diffusion step:
@@
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (2)
where $\boldsymbol{\epsilon}$ is the standard Gaussian noise added during the forward diffusion process. This formulation highlights that at any timestep $t$, the noisy sample $\mathbf{x}_{t}$ is a combination of the original data point $\mathbf{x}_{0}$ and noise $\boldsymbol{\epsilon}$, with their proportions determined by the noise schedule.
Central to diffusion models is the score function $\mathbf{s}(\mathbf{x})=\nabla_{\mathbf{x}}\log p(\mathbf{x})$, which represents the gradient of the log probability density. For data concentrated on a low-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$, the score function has a crucial geometric interpretation: it points approximately normal to the data manifold, with magnitude increasing with distance from the manifold [13, 18]. When training diffusion models, the neural network is typically trained to predict the noise $\boldsymbol{\epsilon}$ that was added during the forward process. This prediction, denoted as $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, aims to approximate the true noise $\boldsymbol{\epsilon}$ used to generate $\mathbf{x}_{t}$ from $\mathbf{x}_{0}$. Importantly, this noise prediction directly relates to the score of the noisy distribution through:
$\mathbf{s}_{t}(\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})=-\frac{\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sigma_{t}}$ (3)
where $\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}}$ is the standard deviation of the noise at timestep $t$. This relationship, derived from the Tweedie-Robbins-Miyasawa formula [19, 20], enables us to extract geometric information about the underlying data manifold without requiring explicit parameterization.
@@ 3.1 Capturing Manifold Geometry via Ambient Metric Deformation: the Stein Score Metric Tensor
"Space tells matter how to move; matter tells spacetime how to curve" C.W. Misner et al. [21]
We define a data-induced metric through geometric deformation [22, 23], drawing inspiration from general relativity, where mass curves spacetime to create gravitational paths. Analogously, our approach deforms the ambient Euclidean space such that geodesics naturally adhere to the data manifold without requiring explicit parameterization (Figure 1 C). Rather than directly characterizing the manifold $\mathcal{M}$, we instead deform the geometry of $\mathbb{R}^{N}$ to accommodate the manifold’s presence—effectively creating a "gravitational pull" toward the data distribution.
@@
[Definition] The Stein score metric tensor is a smooth map $g:\mathbb{R}^{N}\to\text{SPD}(N)$ (the space of symmetric positive-definite $N\times N$ matrices) defined at a point $\boldsymbol{x}\in\mathbb{R}^{N}$ as:
$g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\mathbf{s}(\boldsymbol{x})\mathbf{s}(\boldsymbol{x})^{T},$ (4)
where $\mathbf{I}$ is the $N\times N$ identity matrix, $\mathbf{s}(\boldsymbol{x})=\nabla_{\boldsymbol{x}}\log p(\boldsymbol{x})$ is the score function, and $\lambda$ is a positive penalty parameter that controls the strength of penalization along normal directions.
The Stein score metric $g(\boldsymbol{x})$ equips the ambient space with a Riemannian structure respecting the underlying data manifold geometry. The inner product between vectors $\boldsymbol{u},\boldsymbol{v}\in\mathbb{R}^{N}$ at point $\boldsymbol{x}$ becomes: $<\boldsymbol{u},\boldsymbol{v}>_{g(\boldsymbol{x})}=\boldsymbol{u}^{T}\boldsymbol{v}+\lambda(\mathbf{s}(\boldsymbol{x})^{T}\boldsymbol{u})(\mathbf{s}(\boldsymbol{x})^{T}\boldsymbol{v})$, combining the standard Euclidean inner product with a term that penalizes movement normal to the manifold.
@@
    The metric remains positive definite for all values of $\lambda$ (see Appendix A).
@@ 3.2 Geometric computation on the data manifold
Length of a curve and Energy functionals
A curve in the ambient space $\mathbb{R}^{N}$ is a smooth function $\gamma:[0,1]\rightarrow\mathbb{R}^{N}$. The length of this curve under our score-based metric is computed as:
[…]
@@
[…]
@@
Geodesics
Geodesics are curves with minimal length connecting two points in a Riemannian manifold. In our score-based metric space, geodesics are fundamental as they provide the optimal paths between points while naturally respecting the underlying data manifold structure.
@@
[…]
@@
Geodesic Computation Algorithm
While the geodesic equation provides a mathematical foundation, directly solving this minimization problem in high-dimensional spaces is challenging due to the computational complexity and the irregular behavior of score functions in data-sparse regions. Our approach leverages diffusion models to address these challenges through a three-stage process.
First, we apply controlled noise perturbation by sampling $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ and computing $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}$ for both endpoints. This forward diffusion process serves two purposes: it smooths the optimization landscape by regularizing the energy functional and ensures that identical noise is applied to both endpoints, preserving their relative positioning while making the optimization more stable. The noise level $t$ is a hyperparameter that controls the smoothness of the optimization landscape—higher values provide more stability but potentially less accurate manifold adherence.
With the score function $\mathbf{s}(\mathbf{x})$ obtained from the diffusion model at timestep $t$, we construct our metric tensor (Eq. 4) using a fixed scale parameter $\lambda$. The optimal choice of $\lambda$ depends on both the dataset characteristics and the specific noise level $t$ (see Section 4). Next, we discretize the energy functional and solve it numerically (for more details see B). We represent the path as a sequence of points $\gamma=\{\gamma_{0},\gamma_{1},...,\gamma_{n}\}$, with $\gamma_{0}=\mathbf{x}_{A}$ and $\gamma_{n}=\mathbf{x}_{B}$. The discrete energy becomes:
@@
We minimize this energy using Riemannian gradient descent methods [24] that respect the curved geometry defined by our metric (for more details see Appendix B and Algorithm 2). The gradient computation takes into account how changes in each interior point affect both its contribution to the energy and the score-based metric at that point. We initialize the path with linear interpolation between the endpoints and iteratively refine it until convergence or a maximum iteration count is reached (see Appendix for experiment specific details). Finally, we map the optimized path in noise space back to image space using the denoising capabilities of the diffusion model. Each point along the geodesic is denoised from timestep $t$ back to $t^{\prime}=0$.
@@
Manifold-Aware Interpolation
Armed with our score-based metric tensor and geodesic computation algorithm, we can now perform interpolation between data points that respects the underlying manifold structure. Unlike conventional approaches such as linear interpolation (LERP) [10], spherical interpolation (SLERP) [25, 26, 27] or Noise Diffusion [28] that often cut through low-density regions of the data space, our geodesic-based interpolation follows the natural contours of the data manifold. This quality improvement (see Section 4) stems from the path’s adherence to the manifold structure, which prevents interpolation through low-density "off-manifold" regions where the diffusion model has not been trained.
Manifold-Aware Extrapolation
While interpolation connects two known endpoints, extrapolation extends a path beyond the observed data, requiring a different approach that respects the manifold without a target endpoint. We implement manifold-aware extrapolation through guided walking by the score-metric tensor with momentum.
@@
where $\beta$ controls how strongly the extrapolation maintains its previous direction. We initialize $\mathbf{m}_{0}$ from the tangent direction at the endpoint of the geodesic path (for more details see E). This approach offers several advantages: it maintains coherent progression by preserving directional momentum; it adheres to the manifold through score guidance. With computational complexity of $\mathcal{O}(n\cdot d)$ per step and requiring only one score evaluation per iteration.

### v3

@@ Abstract
Recent advances in diffusion models have demonstrated their remarkable ability to capture complex image distributions, but the geometric properties of the learned data manifold remain poorly understood. We address this gap by introducing a score-based Riemannian metric that leverages the Stein score function from diffusion models to characterize the intrinsic geometry of the data manifold without requiring explicit parameterization. Specifically, our approach defines a metric tensor in the ambient space that stretches distances in directions perpendicular to the manifold while preserving them along tangential directions, effectively creating a geometry where geodesics naturally follow the manifold’s contours. We develop efficient algorithms for computing these geodesics and demonstrate their utility for both interpolation between data points and extrapolation beyond the observed data distribution. Through experiments on synthetic data with known geometry, Rotated MNIST, and complex natural images via Stable Diffusion, we show that our score-based geodesics capture meaningful transformations that respect the underlying data distribution. Our method consistently outperforms baseline approaches on perceptual metrics (LPIPS) and distribution-level metrics (FID, KID), producing smoother, more realistic image transitions. These results reveal the implicit geometric structure learned by diffusion models and provide a principled way to navigate the manifold of natural images through the lens of Riemannian geometry.
@@ 1 Introduction
The geometry of natural images can be studied by viewing them as points on a manifold in high-dimensional ambient – pixel – space. This perspective, commonly known as the manifold hypothesis (Tenenbaum et al., 2000; Bengio et al., 2013), suggests that despite the vast dimensionality of pixel space, natural images concentrate near a much lower-dimensional structure (Carlsson et al., 2008; Henaff et al., 2014). Understanding this geometric structure is crucial for numerous applications including image synthesis, manipulation, and analysis (Goodfellow et al., 2020; Zhu et al., 2017; Mohan et al., 2019), visual perception (DiCarlo & Cox, 2007; Chung et al., 2018), and representation learning (Kingma et al., 2013; Cohen et al., 2020; Chen et al., 2020).
In this work, we propose a data-dependent metric derived from the Stein score function, which can be obtained from a diffusion model trained on a given set of images. Recent advances in diffusion models have demonstrated remarkable capabilities in capturing complex image distributions (Ho et al., 2020; Dhariwal & Nichol, 2021), and several approaches have explored data-dependent metrics to uncover the geometrical properties of image manifolds (Kapusniak et al., 2024; Stanczuk et al., 2024; Wang & Ponce, 2021; Diepeveen et al., 2024; Yun et al., 2024; Park et al., 2023; Samuel et al., 2023). However, these previous methods often require explicit manifold parameterization, are computationally intractable for high-dimensional data, or fail to capture the intrinsic geometric structure that reflects perceptual relationships in the data.
Our approach defines a Riemannian metric in the ambient – pixel – space, leveraging score functions from diffusion models to create a geometry that stretches perpendicular to the data manifold while preserving distances along it. This metric allows us to compute distances and geodesic paths between images directly in pixel space, capturing meaningful relationships that respect the underlying data distribution. Unlike the Euclidean distance, which treats all directions in pixel space equally, our score-based metric accounts for the anisotropic nature of the image manifold, heavily penalizing movement in directions orthogonal to the manifold while preserving natural movement along tangential directions. By generating images along these geodesics, by exploiting a diffusion model, we can visualize and validate these transformations, providing unique insights into the structure of the learned image manifold.
The main contributions of our work are: (i) A novel score-based Riemannian metric that captures the data manifold geometry without explicit parameterization; (ii) Efficient algorithms for computing geodesics and manifold extrapolation in high-dimensional image space; (iii) Empirical validation on both synthetic data with known geometry and complex image data.
@@ 2.1 The Manifold Hypothesis for Natural Images
The manifold hypothesis posits that high-dimensional data, such as natural images, lie approximately on a low-dimensional manifold embedded in the ambient space $\mathbb{R}^{N}$ (Tenenbaum et al., 2000; Bengio et al., 2013; Roweis & Saul, 2000). In this framework, images are represented as points in a $N$-dimensional pixel space, where each axis represents a pixel value. Despite the vast dimensionality of this space – often millions of dimensions for modern high-resolution images –, natural images occupy only a tiny fraction of it due to strong correlations and constraints that restrict the set of plausible pixel configurations (Carlsson et al., 2008; Henaff et al., 2014).
Each set of images can be characterized by its own probability density function $p(\bm{x})$ defined over the pixel space $\mathbb{R}^{N}$. The key insight is that $p(\bm{x})$ concentrates its mass on a much lower-dimensional structure than the ambient dimension $N$, referred to as the support of the data distribution. Under the manifold hypothesis, this support approximates a lower-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$ (Figure 1 A). To capture the geometry of this implicit manifold, we employ data-induced metrics - Riemannian metrics defined in the ambient space that adapt locally according to the underlying probability density (Lee, 1997). Our approach explores how ambient space deformation enables geodesics that naturally follow the manifold’s contours (Figure 1 B,C).
Figure 1: (A) Data points $\mathbf{x}$ sampled from a probability distribution $p(\mathbf{x})$ (left) concentrate on a lower-dimensional manifold $\mathcal{M}$ (right). (B) Linear interpolation (black dashed) versus geodesic interpolation (red curve) between MNIST digits. Geodesics follow the manifold surface, producing valid digit transitions, while linear paths yield superpositions. (C) Geometric deformation (Thorne et al., 2000): data manifold $\mathcal{M}$ embedded in $\mathbb{R}^{N}$ (left) transforms flat Euclidean space into curved metric space $\tilde{\mathbb{R}}^{N}$ (middle, right). In this deformed space, geodesics of $\mathcal{M}$ can be computed directly as geodesics of $\tilde{\mathbb{R}^{N}}$.
@@ 2.2 Diffusion Models & Stein Score
Diffusion models have emerged as powerful generative approaches that gradually transform noise into structured data through an iterative denoising process (Sohl-Dickstein et al., 2015; Ho et al., 2020). These models define a forward process $q$ that progressively adds Gaussian noise to data points $\mathbf{x}_{0}\sim p(\mathbf{x})$ according to a predefined schedule, creating a sequence of increasingly noisy versions $\mathbf{x}_{t}$ with $t$ representing the diffusion step:
@@
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (2)
where $\bm{\epsilon}$ is the standard Gaussian noise added during the forward diffusion process. This formulation highlights that at any timestep $t$, the noisy sample $\mathbf{x}_{t}$ is a combination of the original data point $\mathbf{x}_{0}$ and noise $\bm{\epsilon}$, with their proportions determined by the noise schedule.
The score function $\mathbf{s}(\mathbf{x})=\nabla_{\mathbf{x}}\log p(\mathbf{x})$ is central to diffusion models, which represents the gradient of the log probability density. For data concentrated on a low-dimensional manifold $\mathcal{M}\subset\mathbb{R}^{N}$, the score function has a crucial geometric interpretation: it points approximately normal to the data manifold, with magnitude increasing with distance from the manifold (Stanczuk et al., 2024; Pidstrigach, 2022; Yun et al., 2024). When training diffusion models, the neural network is typically trained to predict the noise $\bm{\epsilon}$ that was added during the forward process. This prediction, denoted as $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, aims to approximate the true noise $\bm{\epsilon}$ used to generate $\mathbf{x}_{t}$ from $\mathbf{x}_{0}$. Importantly, this noise prediction directly relates to the score of the noisy distribution through:
$\mathbf{s}_{t}(\mathbf{x}_{t})=\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})=-\frac{\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)}{\sigma_{t}}$ (3)
where $\sigma_{t}=\sqrt{1-\bar{\alpha}_{t}}$ is the standard deviation of the noise at timestep $t$. This relationship, derived from the Tweedie-Robbins-Miyasawa formula (Robbins, 1956; Miyasawa et al., 1961), enables us to extract geometric information about the underlying data manifold without requiring explicit parameterization.
@@ 3.1 Capturing Manifold Geometry via Ambient Metric Deformation: the Stein Score Metric Tensor
”Space tells matter how to move; matter tells spacetime how to curve” C.W. Misner et al. (Thorne et al., 2000)
We define a data-induced metric through geometric deformation (Simon, 2002; Ehrlich, 1974), drawing inspiration from general relativity, where mass curves spacetime to create gravitational paths. Analogously, our approach deforms the ambient Euclidean space such that geodesics naturally adhere to the data manifold without requiring explicit parameterization (Figure 1 C). Rather than directly characterizing the manifold $\mathcal{M}$, we instead deform the geometry of $\mathbb{R}^{N}$ to accommodate the manifold’s presence—effectively creating a ”gravitational pull” toward the data distribution.
@@
[Definition] The Stein score metric tensor is a smooth map $g:\mathbb{R}^{N}\to\text{SPD}(N)$ (the space of symmetric positive-definite $N\times N$ matrices) defined at a point $\bm{x}\in\mathbb{R}^{N}$ as:
$g(\bm{x})=\mathbf{I}+\lambda\cdot\mathbf{s}(\bm{x})\mathbf{s}(\bm{x})^{T},$ (4)
where $\mathbf{I}$ is the $N\times N$ identity matrix, $\mathbf{s}(\bm{x})=\nabla_{\bm{x}}\log p(\bm{x})$ is the score function, and $\lambda$ is a positive penalty parameter that controls the strength of penalization along normal directions (see Appendix D.2 for discussion on optimal $\lambda$).
The Stein score metric $g(\bm{x})$ equips the ambient space with a Riemannian structure respecting the underlying data manifold geometry. The inner product between vectors $\bm{u},\bm{v}\in\mathbb{R}^{N}$ at point $\bm{x}$ becomes: $\bm{u},\bm{v}>_{g(\bm{x})}=\bm{u}^{T}\bm{v}+\lambda(\mathbf{s}(\bm{x})^{T}\bm{u})(\mathbf{s}(\bm{x})^{T}\bm{v})$, combining the standard Euclidean inner product with a term that penalizes movement normal to the manifold.
@@
    The metric remains positive definite for all values of $\lambda$ (see Appendix B).
Note that our metric construction finds theoretical grounding in the framework of rank-1 metric perturbations introduced by  Hartmann et al. (2022), who demonstrated that metrics of the form $G(x)=I+\lambda\nabla f(x)\nabla f(x)^{T}$, where $f(x)=\log p(x)$ create anisotropic Riemannian structures that preferentially penalize movement along gradient directions (see Appendix A).
@@ 3.2 Geometric computation on the data manifold
Length of a curve and Energy functionals. A curve in the ambient space $\mathbb{R}^{N}$ is a smooth function $\gamma:[0,1]\rightarrow\mathbb{R}^{N}$. The length of this curve under our score-based metric is computed as:
$\mathcal{L}[\gamma]=\int_{0}^{1}\|\dot{\gamma}(\tau)\|_{g(\gamma(\tau))}d\tau=\int_{0}^{1}\sqrt{\dot{\gamma}(\tau)^{T}\dot{\gamma}(\tau)+\lambda(\mathbf{s}(\gamma(\tau))^{T}\dot{\gamma}(\tau))^{2}}\,d\tau,$ (5)
@@
$\mathcal{E}[\gamma]=\frac{1}{2}\int_{0}^{1}\|\dot{\gamma}(\tau)\|^{2}_{g(\gamma(\tau))}d\tau=\frac{1}{2}\int_{0}^{1}\left[\|\dot{\gamma}(\tau)\|^{2}+\lambda(\mathbf{s}(\gamma(\tau))^{T}\dot{\gamma}(\tau))^{2}\right]d\tau.$ (6)
@@
Geodesics. Geodesics are curves with minimal length connecting two points in a Riemannian manifold. In our score-based metric space, geodesics are fundamental as they provide the optimal paths between points while naturally respecting the underlying data manifold structure.
@@
$\gamma^{*}=\arg\min_{\begin{subarray}{c}\gamma\\ \gamma(0)=\mathbf{x}_{A},\gamma(1)=\mathbf{x}_{B}\end{subarray}}\mathcal{E}[\gamma]=\arg\min_{\begin{subarray}{c}\gamma\\ \gamma(0)=\mathbf{x}_{A},\gamma(1)=\mathbf{x}_{B}\end{subarray}}\frac{1}{2}\int_{0}^{1}\|\dot{\gamma}(\tau)\|^{2}_{g(\gamma(\tau))}d\tau,$ (7)
@@
Geodesic Computation Algorithm. While the geodesic equation provides a mathematical foundation, directly solving this minimization problem in high-dimensional spaces is challenging due to the computational complexity and the irregular behavior of score functions in data-sparse regions. Our approach leverages diffusion models to address these challenges through a three-stage process.
First, we apply controlled noise perturbation by sampling $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ and computing $\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$ for both endpoints. This forward diffusion process serves two purposes: it smooths the optimization landscape by regularizing the energy functional and ensures that identical noise is applied to both endpoints, preserving their relative positioning while making the optimization more stable. The noise level $t$ is a hyperparameter that controls the smoothness of the optimization landscape—higher values provide more stability but potentially less accurate manifold adherence.
With the score function $\mathbf{s}(\mathbf{x})$ obtained from the diffusion model at timestep $t$, we construct our metric tensor (Eq. 4) using a fixed scale parameter $\lambda$. The optimal choice of $\lambda$ depends on both the dataset characteristics and the specific noise level $t$ (see Section 4). Next, we discretize the energy functional and solve it numerically (for more details see E). We represent the path as a sequence of points $\gamma=\{\gamma_{0},\gamma_{1},...,\gamma_{n}\}$, with $\gamma_{0}=\mathbf{x}_{A}$ and $\gamma_{n}=\mathbf{x}_{B}$. The discrete energy becomes:
@@
We minimize this energy using Riemannian gradient descent methods (Bonnabel, 2013) that respect the curved geometry defined by our metric (for more details see Appendix E and Algorithm 2). The gradient computation takes into account how changes in each interior point affect both its contribution to the energy and the score-based metric at that point. We initialize the path with linear interpolation between the endpoints and iteratively refine it until convergence or a maximum iteration count is reached (see Appendix for experiment specific details). Finally, we map the optimized path in noise space back to image space using the denoising capabilities of the diffusion model. Each point along the geodesic is denoised from timestep $t$ back to $t^{\prime}=0$.
@@
Manifold-Aware Interpolation. Armed with our score-based metric tensor and geodesic computation algorithm, we can now perform interpolation between data points that respects the underlying manifold structure. Unlike conventional approaches such as linear interpolation (LERP) (Ho et al., 2020), spherical interpolation (SLERP) (Shoemake, 1985; Song et al., 2020a; b) or Noise Diffusion (Zheng et al., 2024) that often cut through low-density regions of the data space, our geodesic-based interpolation follows the natural contours of the data manifold. This quality improvement (see Section 4) stems from the path’s adherence to the manifold structure, which prevents interpolation through low-density ”off-manifold” regions where the diffusion model has not been trained.
Manifold-Aware Extrapolation. While interpolation connects two known endpoints, extrapolation extends a path beyond the observed data, requiring a different approach that respects the manifold without a target endpoint. We implement manifold-aware extrapolation through guided walking by the score-metric tensor with momentum.
@@
where $\beta$ controls how strongly the extrapolation maintains its previous direction. We initialize $\mathbf{m}_{0}$ from the tangent direction at the endpoint of the geodesic path (for more details see H). This approach offers several advantages: it maintains coherent progression by preserving directional momentum; it adheres to the manifold through score guidance. With computational complexity of $\mathcal{O}(n\cdot d)$ per step and requiring only one score evaluation per iteration.